write an equation for the following function in x-intercept form:
it shows the parabola,
and the following points are labeled:
(3,0) (7,0)
i think they're intercepts?
and then (6, -9)

Answer by stanbon(75887)

write an equation for the following function in x-intercept form:
it shows the parabola,
and the following points are labeled:
(3,0) (7,0)
i think they're intercepts?
---
Yes, they are roots, so the function has (x-3) and (x-7) as factors.
------------------
and then (6, -9)
-------------------------
Equation:
y = a(x-3)(x-7)
----
Since the function passes thru (6,-9) substitute to solve for "a":
-9 = a(6-3)(6-7)
-9 = a*3*-1
a = -9/-3 = 3
----------------------
Equation: y = 3(x-3)(x-7)
==================================
Cheers,
Stan H.